### 11 Apr 2026 v2

# Point modules over the universal enveloping algebras of color Lie algebras

Abstract.
Let $k$ be an algebraically closed field with characteristic 0. In this paper, we define the notion of a $q^{\prime}$-Heisenberg normal element of a $\mathbb{Z}$-graded $k$-algebra. This $q^{\prime}$-Heisenberg normal element gives the structure of some sets of modules related to point modules. Also, we determine the set of point modules over an Artin–Schelter regular algebra obtained as the universal enveloping algebra of a color Lie algebra. Moreover, we give a concrete integer such that the inverse system of truncated point schemes of it is constant.
Key words and phrases:
AS-regular algebras, point modules, point schemes, color Lie algebras
2020 Mathematics Subject Classification:
Primary 14A22; Secondary 16S38, 17B35, 17B75.
1. Introduction
[…]
An Artin–Schelter regular algebra (or an AS-regular algebra, for shortly) was introduced by Artin–Schelter [1]. By the fact that a commutative AS-regular algebra is a polynomial algebra, an AS-regular algebra is regarded as a “noncommutative polynomial algebra”. In the rest of this paper, we assume that an AS-regular algebra is generated in degree $1$. An AS-regular algebra was often studied by a geometric method. One of the geometric methods is studying a point scheme introduced by Artin–Tate–Van den Bergh [2]: Let $A=k\langle x_{0},x_{1},\cdots,x_{s}\rangle/\langle f_{1},\cdots,f_{t}\rangle$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, $V$ the $k$-vector space $\langle x_{0},x_{1},\cdots,x_{s}\rangle_{k}$, and $I$ the homogeneous two-sided ideal $\langle f_{1},\cdots,f_{t}\rangle$ of $k\langle V\rangle$. For an integer $i\geq 1$, an element of $I_{i}$ defines a multilinear function on $(V^{*})^{\times i}$. The $i$-th truncated point scheme $\Gamma_{i}(A)\subset(\mathbb{P}^{s})^{\times i}$ of $A$ is defined as the zero scheme defined by $I_{i}$. Also, $\mathrm{pr}^{(i+1)}_{1,i}:\Gamma_{i+1}(A)\rightarrow\Gamma_{i}(A)$ is the restriction of the projection $(\mathbb{P}^{s})^{\times i+1}\rightarrow(\mathbb{P}^{s})^{\times i}$ of the first $i$ elements. The point scheme $\Gamma(A)$ of $A$ is defined as the inverse limit of $\Gamma_{i}(A)$ : $\Gamma(A):=\displaystyle{\lim_{\leftarrow}\Gamma_{i}(A)}$ ([2]).
For a $3$-dimensional AS-regular algebra $A$, the point scheme $\Gamma(A)$ of $A$ is isomorphic to a product of the projective space either $\mathbb{P}^{2}$ or $\mathbb{P}^{1}\times\mathbb{P}^{1}$, or a divisor of either. In [2], $3$-dimensional AS-regular algebras were classified by point schemes. However, for a $4$ or more dimensional AS-regular algebra, the dimension of the (truncated) point scheme can be low. For $4$-dimensional AS-regular algebras with $3$ or more Gelfand–Kirillov dimensions, Lu–Palmieri–Wu–Zhang [16] gives the list of the number of generators and the number of relations with degrees. Let $R$ be a set of relations which have degrees on this list, then the dimension of a truncated point scheme of $R$ is expected to be zero from Chirvasitu–Kanda [9]. Moreover, by Vancliff–Van Rompay–Willaert [27] and Vancliff–Veerapen [28], some concrete examples were constructed in which the $4$-dimensional AS-regular algebras have the 0-dimensional point scheme. Also, in Vancliff [29], a $5$-dimensional AS-regular algebra without any closed points in the point scheme was constructed. On the other hand, the point variety of a skew polynomial algebra, which is a quadratic AS-regular algebra, can have the “large” point scheme (Belmans–De Laet–Le Bruyn[6] and Vitoria[30]). In this paper, by using color Lie algebras, we construct some dimensional AS-regular algebras with the “large” point scheme and relations having some degrees.
A color Lie algebra was introduced by Ree [22] as a generalized Lie algebra. A one of noncommutative algebraic geometrical studies in an early stage, it was studied as a schematic algebra by Van Oystaeyen–Willaert [26]. Also, for a color Lie algebra $L$, the universal enveloping algebra of $L$ is an AS-regular algebra under some conditions.
A module called a point module over a connected $\mathbb{N}$-graded finitely presented $k$-algebra AA generated in degree $1$ is important from the following fact ([2]): The set of closed points of the point scheme of $A$ corresponds one-to-one to the set of isomorphism classes of point modules over $A$. In this paper, in order to study a point module, we define the notion of a $q^{\prime}$-Heisenberg normal element of a $\mathbb{Z}$-graded $k$-algebra.
Definition (Definition 3.1).
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra and $g$ a regular homogeneous normal element of $A$ with degree $n\geq 1$. A regular homogeneous normal element $g$ is called a $q^{\prime}$-Heisenberg normal element of $A$, if there exist $u\in k^{\times}$, $x\in A_{1}$, and $y\in A_{n-1}$ satisfies the following equations;
  1. (i)
    g=xyuyxg=xy-uyx;
  2. (ii)
    0=xgugx0=xg-ugx;
  3. (iii)
    0=gyuyg0=gy-uyg.
The relations in the above definition of a $q^{\prime}$-Heisenberg normal element are similar to defining relations of another $q$-analog of the universal enveloping algebra of the Heisenberg algebra (cf. Kirkman–Small [13]). To emphasize a correspondence to another case, we denote $q^{\prime}$ instead of simply denoting $q$. This $q^{\prime}$-Heisenberg normal element gives us the structure of some sets of modules having a relationship to a point module. Assume that $A$ is a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and $g$ is a $q^{\prime}$-Heisenberg normal element of $A$. Proposition 3.7 shows that the set of base-point modules over $A$ completely overlaps one of $A/\langle g\rangle$. Also, in Theorem 4.3, we obtain a criterion for the nonexistence of a truncated point module. Therefore, we obtain the fact that some set of truncated point modules over $A$ completely overlaps one of $A/\langle g\rangle$. In Theorem 4.5, using this result, we determine the structure of the set of closed points of the point scheme of the universal enveloping algebra of color Lie algebras.
For a connected $\mathbb{N}$-graded finitely presented $k$-algebra AA, Artin, Tate, and Van den Bergh raised the question [2, Question 3.16] as follows: When $A$ is noetherian, is the inverse system $\{\Gamma_{i}(A)\}$ constant for large $i$? A partial answer to this question was given by Artin and Zhang [4]: When $A$ is strongly noetherian, $\{\Gamma_{i}(A)\}$ is constant for some $i>0$. In Corollary 4.6, for the universal enveloping algebra $U(L)$ of a color Lie algebra $L$, we give a concrete integer $i$ such that $\{\Gamma_{i}(U(L))\}$ is constant. Note that this integer $i$ is obtained by a property of the color Lie algebra.
This paper is organized as follows: In Section 2, we recall some definitions and properties of the quasi-Veronese algebra defined by Mori [18], the twisted algebra defined by Zhang [31], the dehomogenization, point modules, and color Lie algebras. In Section 3, we define the notion of a $q^{\prime}$-Heisenberg normal element and prove basic properties. In Proposition 3.7, we prove that the non-existence of a $g$-torsionfree base-point module for a $q^{\prime}$-Heisenberg normal element $g$. In Section 4, we prove Theorem 4.3, which is a theorem of the nonexistence of a $g$-torsionfree module with some conditions. By Theorem 4.3, we determine the set of closed points of the point schemes of the universal enveloping algebra of color Lie algebras in Theorem 4.5. At the end of this section, we show that the inverse system of the truncated point schemes of it is constant in Corollary 4.6.
2. Preliminaries
2.1. Graded algebras and filtered algebras
In this subsection, we recall the categorical equivalences obtained by the quasi-Veronese algebra [18], the twisted algebra [31], and the dehomogenization (see [15]). First, we recall the definition of the quasi-Veronese algebra.
Definition 2.1 ([18, Definition 3.7]).
Let $A=\bigoplus_{i\in\mathbb{Z}}A_{i}$ be a $\mathbb{Z}$-graded $k$-algebra and $r\geq 1$ a positive integer. A $\mathbb{Z}$-graded $k$-algebra $A^{[r]}$ called the $r$-th quasi-Veronese algebra of $A$ is defined as follows:
$A^{[r]}:=\bigoplus_{i\in\mathbb{Z}}\left(\begin{array}[]{cccc}A_{ri}&A_{ri+1}&\cdots&A_{ri+r-1}\\ A_{ri-1}&A_{ri}&\cdots&A_{ri+r-2}\\ \vdots&&\ddots&\vdots\\ A_{ri-r+1}&A_{ri-r+2}&\cdots&A_{ri}\\ \end{array}\right)$
with the multiplication $(a_{i,j})(b_{i,j})=\left(\sum_{l=0}^{r-1}a_{l,j}b_{i,l}\right)$ for $(a_{i,j})\in(A^{[r]})_{p}$, $(b_{i,j})\in(A^{[r]})_{q}$ for $p,q\in\mathbb{Z}$.
The definition of the quasi-Veronese algebra gives the following categorical equivalence:
Lemma 2.2 ([18, Lemma 3.9]).
For a positive integer $r\geq 1$, let $A=\bigoplus_{i\in\mathbb{Z}}A_{i}$ be a $\mathbb{Z}$-graded $k$-algebra and $A^{[r]}$ the $r$-th quasi-Veronese algebra of $A$. Then the following categorical equivalence holds:
[…]
$Q(M):=\bigoplus_{i\in\mathbb{Z}}\left(\bigoplus_{j=0}^{r-1}M_{ir-j}\right)$
with the action $(a_{i,j})(m_{i})=\left(\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{l=0}^{r-1}a_{l,i}m_{l}\right)$ for $(a_{i,j})\in(A^{[r]})_{p},(m_{i})\in(Q(M))_{q}$ for $p,q\in\mathbb{Z}$.
Next, we recall the definition of the twisted algebra due to Zhang [31]. Here, to deal with the left module, we recall the left version of it.
Definition 2.3 ([31, Definition 4.1]).
Let $A=\bigoplus_{i\in\mathbb{Z}}A_{i}$ be a $\mathbb{Z}$-graded $k$-algebra. A set of graded $k$-linear automorphisms $\nu=\{\nu_{i}\mid i\in\mathbb{Z}\}$ is called a ($\ell$-)twisting system if $\nu$ satisfies the following equation:
$\nu_{l}(\nu_{j}(a)b)=\nu_{j+l}(a)\nu_{l}(b)\,\text{ for }i,j,l\in\mathbb{Z},a\in A_{i},\text{ and }b\in A_{j}.$
Definition 2.4 ([31, Proposition and Definition 4.2]).
Let $A=\bigoplus_{i\in\mathbb{Z}}A_{i}$ be a $\mathbb{Z}$-graded $k$-algebra and $\nu=\{\nu_{i}\mid i\in\mathbb{Z}\}$ a ($\ell$-)twisting system of $A$. A $\mathbb{Z}$-graded $k$-algebra $^{\nu}A$ called the ($\ell$-)twisted algebra is defined as follows:
$^{\nu}A:=\operatorname*{\mathchoice{\Xdbigoplus}{\Xdbigoplus}{\Xtbigoplus}{\Xtbigoplus}}\limits_{i\in\mathbb{Z}}A_{i}$ with the multiplication $a\circ b:=\nu_{j}(a)b$ where $a\in{}^{\nu}A_{i}$, $b\in{}^{\nu}A_{j}$, and $i,j\in\mathbb{Z}$.
The definition of the ($\ell$-)-twisted algebra gives the equivalence of the (left) graded module categories.
Proposition 2.5 ([31, Corollary 4.4]).
Let $A=\bigoplus_{i\in\mathbb{Z}}A_{i}$ be a $\mathbb{Z}$-graded $k$-algebra and $\nu$ a ($\ell$-)twisting system of $A$. Then the following categorical equivalence holds:
[…]
$^{\nu}M:=\bigoplus_{i\in\mathbb{Z}}M_{i}$ with the action $a\circ m:=\nu_{j}(a)m$ where $a\in{}^{\nu}A_{i}$, $m\in{}^{\nu}M_{j}$, and $i,j\in\mathbb{Z}$.
We recall the definition and some facts of the dehomogenization (see [15] for details).
Definition 2.6 (see [15, Chapter I. 4.3]).
Let $A=\bigoplus_{i\in\mathbb{Z}}A_{i}$ be a $\mathbb{Z}$-graded $k$-algebra and $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ a graded $A$-module, $T$ a regular central element of $A$ with degree $1$. A filtered ring $\mathcal{R}$ with the filtration $F\mathcal{R}=\{F_{i}\mathcal{R}\}_{i\in\mathbb{Z}}$ called the dehomogenization of $A$ with respect to $T$ is defined as
$\mathcal{R}:=A/\langle 1-T\rangle,\quad F_{i}\mathcal{R}:=(A_{i}+\langle 1-T\rangle)/\langle 1-T\rangle\text{ for }i\in\mathbb{Z}.$
Also, a filtered $R$-module $\mathcal{M}$ with the filtration $F\mathcal{M}=\{F_{i}\mathcal{M}\}_{i\in\mathbb{Z}}$ called the dehomogenization of $M$ with respect to $T$ is defined as follows:
$\mathcal{M}:=M/\langle 1-T\rangle M,\quad F_{i}\mathcal{M}:=(M_{i}+\langle 1-T\rangle M)/\langle 1-T\rangle M\text{ for }i\in\mathbb{Z}.$
The dehomogenization satisfies some nice properties as stated below
Lemma 2.7 (see [15, Chapter I 4.3 Lemma 2]).
Let $A=\bigoplus_{i\in\mathbb{Z}}A_{i}$ be a $\mathbb{Z}$-graded $k$-algebra, $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ a graded $A$-module, and $T$ a regular central element of $A$ with degree $1$. Then the equation $(1-T)A\cap A_{i}=0$ holds for $i\in\mathbb{Z}$. Also, if $M$ is a $T$-torsionfree module, then the equation $(1-T)M\cap M_{i}=0$ holds for $i\in\mathbb{Z}$.
Proposition 2.8 (See [15, Chapter I. 4.3.6]).
[…]
2.2. Point schemes and point modules
Let $A=k\langle x_{0},x_{1},\cdots,x_{s}\rangle/\langle f_{1},\cdots,f_{t}\rangle$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ where generators $x_{i}$ with degree $1$ and relations $f_{j}$ with degree $2$ or more. In this subsection, we recall the definition and the propositions of a point module of $A$.
Definition 2.9 ([2, 19]).
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$.
  1. (1)
    A graded module $P=\bigoplus_{i\in\mathbb{Z}}P_{i}$ is called a point module over $A$ if $P=AP_{0}$ and $P$ satisfies the equation; $\operatorname{dim}_{k}P_{i}=\begin{cases}1&\text{for $i\geq 0$},\\ 0&\text{for $i<0$}.\end{cases}$
  2. (2)
    For each integer $d\geq 1$, a graded module $P=\bigoplus_{i\in\mathbb{Z}}P_{i}$ is called a truncated point module of length $d+1$ over $A$ if $P=AP_{0}$ and $P$ satisfies the equation; $\operatorname{dim}_{k}P_{i}=\begin{cases}1&\text{for $0\leq i\leq d$},\\ 0&\text{otherwise}.\end{cases}$
Moreover, we recall the definition of a base-point module.
Definition 2.10 ([7]).
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and $c\geq 1$ an integer. A graded module $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ is called a base-point module over $A$ if $AM_{0}, MM$ satisfies the equation;
$\operatorname{dim}_{k}M_{i}=\begin{cases}c&\text{for $i\geq 0$};\\ 0&\text{for $i<0$};\end{cases}$
and $M$ is $1$-critical with respect to the Gelfand–Kirillov dimension.
Remark that, for a base-point module $M$, when $c=1$, $M$ is a point module. Also, when $c\geq 2$, $M$ is called a fat point module ([5]).
For a $k$-algebra AA, an element $g$ of $A$ satisfying an equation $gA=Ag$ is called a normal element. Also, for an $A$-module $M$, $M$ is called a $g$-torsionfree module if $gm=0$ implies $m=0$ for $m\in M$. To recall the well-known property for a point module and a normal element, we show a useful property as follows:
Lemma 2.11.
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and $P=\bigoplus_{i=0}^{d}km_{i}$ a truncated point module of length $d+1$ over $A$ (resp. $P^{\prime}=\bigoplus_{j\in\mathbb{Z}}km^{\prime}_{j}$ is a point module over $A$). Suppose that $g$ is a homogeneous normal element of $A$ with degree $n$. Then $gm_{i-1}=0$ holds if and only if $gm_{i}=0$ holds for $1\leq i\leq d-n$ (resp. $gm^{\prime}_{j-1}=0$ holds if and only if $gm^{\prime}_{j}=0$ holds for $j\geq 1$).
Proof.
Since a truncated point module $P$ over $A$ is cyclic, there exists an element of $a\in A_{1}$ such that $am_{i-1}=m_{i}$. Let $a^{\prime}$ be an element satisfying $a^{\prime}g=ga$. If $gm_{i-1}=0$, then $gm_{i}=gam_{i-1}=a^{\prime}gm_{i-1}=0$. Conversely, we assume that $gm_{i}=0$ and $gm_{i-1}\neq 0$. Let $b,b^{\prime}\in A_{1}$ be elements satisfying $bgm_{i-1}=m_{i+n}$ and $bg=gb^{\prime}$. The equation $m_{i+n}=bgm_{i-1}=gb^{\prime}m_{i-1}=0$ leads to a contradiction. ∎
By Lemma 2.11, a useful property corresponding to a $g$-torsionfree property of a module is defined as in Definition 2.12.
Definition 2.12.
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and $g$ a homogeneous normal element of $A$ with degree $n$. For an integer $d\geq n$, a truncated point module $P=\bigoplus_{i=0}^{d}P_{i}$ of length $d+1$ over $A$ is called a truncated $g$-torsionfree point module of length $d+1$ over $A$ if the following equation satisfies:
$0=\left\{a\in\operatorname*{\mathchoice{\Xdbigoplus}{\Xdbigoplus}{\Xtbigoplus}{\Xtbigoplus}}\limits_{i=0}^{d-n}P_{i}\middle|ga=0\right\}.$
Proposition 2.13 is a generalized one of well-known properties of a point module over $A$.
Proposition 2.13.
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and $P$ a truncated point module of length $d+1$ over $A$ (resp. $P^{\prime}$ a point module over $A$). If $P$ is not a truncated $g$-torsionfree point module of length $d+1$ over $A$ (resp. $P^{\prime}$ is not a $g$-torsionfree point module over $A$), then $gP=0$ (resp. $gP^{\prime}=0$).
Proof.
By Lemma 2.11. ∎
For a connected $\mathbb{N}$-graded finitely presented $k$-algebra AA generated in degree $1$, Artin, Tate, and Van den Bergh [2] proved a nice relationship between the point scheme $\Gamma(A)$ of $A$ and the isomorphism classes of point modules over $A$.
Proposition 2.14 ([2, Proposition 3.9, Corollary 3.13]).
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and $\Gamma_{d}(A)$ the $d$-th truncated point scheme of $A$ (resp. $\Gamma(A)$ the point scheme of $A$). Then there exists a one-to-one correspondence between the set of isomorphism classes of truncated point modules of length $d+1$ over $A$ (resp. point modules over $A$) and the set of closed points of $\Gamma_{d}(A)$ (resp. $\Gamma(A)$).
2.3. Color Lie algebras
In this subsection, we recall the definitions and properties of the universal enveloping algebra of a color Lie algebra. First, we recall the definitions of a skew symmetric bicharacter and a color Lie algebra.
Definition 2.15 ([22]).
Let $G$ be an abelian group. A map $\varepsilon:G\times G\rightarrow k^{\times}$ is a skew symmetric bicharacter if, for $\alpha,\beta,\gamma\in G$, the following equations hold;
  1. (i)
    $\varepsilon(\alpha,\beta)\varepsilon(\beta,\alpha)=1$;
  2. (ii)
    $\varepsilon(\alpha+\beta,\gamma)=\varepsilon(\alpha,\gamma)\varepsilon(\beta,\gamma)$;
  3. (iii)
    $\varepsilon(\alpha,\beta+\gamma)=\varepsilon(\alpha,\beta)\varepsilon(\alpha,\gamma)$.
Definition 2.16 ([22]).
Let $G$ be an abelian group and $\varepsilon:G\times G\rightarrow k^{\times}$ is a skew symmetric bicharacter. Suppose that $L=\bigoplus_{\gamma\in G}L_{\gamma}$ is a $G$-graded $k$-vector space and $L\times L\rightarrow L$ is a $G$-graded bilinear form. A $(G,\varepsilon)$-color Lie algebra is the pair L=(L,[,])L=(L,[\,\,,\,\,]) satisfying equations as follows for $a,b,c\in h(L)$;
  1. (i)
    $a,b]=-\varepsilon(|a|,|b|)[b,a]$;
  2. (ii)
    $\varepsilon(|c|,|a|)[a,[b,c]]+\varepsilon(|a|,|b|)[b,[c,a]]+\varepsilon(|b|,|c|)[c,[a,b]]=0$.
Additionally, the subspace of $L$ is defined as $L_{-}:=\bigoplus_{\gamma\in G_{-}}L_{\gamma}$ (resp. $L_{+}:=\bigoplus_{\gamma\in G_{+}}L_{\gamma}$) with $G_{+}:=\{\gamma\in G\mid\varepsilon(\gamma,\gamma)=+1\}$ (resp. $G_{-}:=\{\gamma\in G\mid\varepsilon(\gamma,\gamma)=-1\}$).
Next, for a $(G,\varepsilon)$-color Lie algebra $L$, we recall the definition of the universal enveloping algebra of $L$ and the $\varepsilon$-symmetric algebra of $L$.
Definition 2.17 ([22]).
Let $L$ be a $(G,\varepsilon)$-color Lie algebra, $T(L)$ the tensor algebra of $L$, and $J(L)$ the two-sided ideal of $T(L)$ generated by the relations $a\otimes b-\varepsilon(|a|,|b|)b\otimes a-\langle a,b\rangle$ for any $a,b\in h(L)$. The universal enveloping algebra of $L$ is defined as
$U(L):=T(L)/J(L).$
The universal enveloping algebra $U(L)$ of a $(G,\varepsilon)$-color Lie algebra $L$ has natural $G$-gradtion; the degree of $x_{1}\otimes\cdots\otimes x_{m}$ is $|x_{1}|+\cdots+|x_{m}|$ where $x_{i}\in h(L)$ for $1\leq i\leq m$.
Also, for a $(G,\varepsilon)$-color Lie algebra $L$, the $\varepsilon$-symmetric algebra $S_{\varepsilon}(L)$ of $L$ as the quotient of $T(L)$ by the two-sided ideal generated by the relations:
$a\otimes b-\varepsilon(|a|,|b|)b\otimes a\text{ for }a,b\in h(L).$
For more details, see [24]. Note that $\varepsilon$-symmetric algebra can be defined for a $G$-graded $k$-vector space and a skew symmetric bicharacter $\varepsilon:G\times G\rightarrow k^{\times}$.
Remark 2.18.
  1. (1)
    The universal enveloping algebra $U(L)$ of a finite dimensional color Lie algebra $L$ with $L_{-}=0$ has a finite Gelfand–Kirillov dimension by [17, 8.1.14] and [24, 4.C] (also, see [22]).
  2. (2)
    For the universal enveloping algebra $U(L)$ of a finite dimensional $(\mathbb{Z}^{m},\varepsilon)$-color Lie algebra $L$ with $L_{-}=0$, $U(L)$ is regarded as $\mathbb{Z}$-graded $k$-algebra by the following way;
    $U(L)=\bigoplus_{i\in\mathbb{Z}}\left(\bigoplus_{i=a_{1}+\cdots a_{m}}U(L)_{(a_{1},\cdots,a_{m})}\right).$
    In particular, when a $\mathbb{Z}$-graded $k$-algebra $U(L)$ is a connected $\mathbb{N}$-graded $k$-algebra generated in degree $1$, $U(L)$ is an AS-regular algebra by [14, 6.3 Theorem] and [20, Theorems 3.1, 3.2].
In the end of this subsection, for the universal enveloping algebra $U(L)$ of a $(G,\varepsilon)$-color Lie algebra $L$, we recall a $G$-graded free resolution of the trivial module $k$ of $U(L)$ called the color Koszul resolution of $U(L)$ ([8, Theorem 3]).
Theorem 2.19.
Let L=(V,[,])L=(V,[\,,\,]) be a $(G,\varepsilon)$-color Lie algebra and $U(L)$ the universal enveloping algebra of $L$. For $r\geq 0$, the free module $C_{r}$ is defined as $C_{r}=U(L)\otimes_{k}\bigwedge_{\varepsilon}^{r}V$ where $\bigwedge_{\varepsilon}^{r}V$ is the $r$-th part of $\bigwedge_{\varepsilon}=T(V)/\langle u\otimes v+\varepsilon(|u|,|v|)v\otimes u\mid u,v\in h(V)\rangle$. Also, for $r\geq 1$, $d_{r}:C_{r}\rightarrow C_{r-1}$ is defined as follows:
$a\otimes$ $(v_{1}\wedge\cdots\wedge v_{r})$
$\mapsto$ $\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{i=1}^{r}(-1)^{i+1}\eta_{i}av_{i}\otimes\left(v_{1}\wedge\cdots\wedge\widehat{v_{i}}\wedge\cdots\wedge v_{r}\right)$
$+\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{1\leq i<j\leq r}(-1)^{i+j}\eta_{i}\eta_{j}\varepsilon(|v_{j}|,|v_{i}|)a\otimes\left([v_{i},v_{j}]\wedge v_{1}\wedge\cdots\wedge\widehat{v_{i}}\wedge\cdots\wedge\widehat{v_{j}}\wedge\cdots\wedge v_{r}\right)$
where $v_{i}\in h(L)$, $a\in h(U(L))$, $\eta_{1}=1$, and $\eta_{i}=\Pi_{l=1}^{i-1}\varepsilon(|v_{l}|,|v_{i}|)$. The sign $\widehat{}$ indicates that the element below it must be omitted. Then the sequence
$\cdots\rightarrow C_{i}\xrightarrow{d_{i}}C_{i-1}\rightarrow\cdots\rightarrow C_{2}\xrightarrow{d_{2}}C_{1}\xrightarrow{d_{1}}C_{0}\rightarrow k\rightarrow 0$
is a $G$-graded free resolution of the trivial module $k$ as a $G$-graded $U(L)$-module.
3. $q^{\prime}$-Heisenberg normal elements
In this section, for a $\mathbb{Z}$-graded $k$-algebra AA, we define a $q^{\prime}$-Heisenberg normal element of $A$ and prove some basic properties about this.
Definition 3.1.
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra and $g$ a regular homogeneous normal element of $A$ with degree $n\geq 1$. A regular homogeneous normal element $g$ is called a $q^{\prime}$-Heisenberg normal element of $A$, if there exist $u\in k^{\times}$, $x\in A_{1}$, and $y\in A_{n-1}$ satisfies the following equations;
  1. (i)
    g=xyuyxg=xy-uyx;
  2. (ii)
    0=xgugx0=xg-ugx;
  3. (iii)
    0=gyuyg0=gy-uyg.
Remark 3.2.
If either $x$ or $y$ is a central element of $A$, then $g=0$ holds. In particular, when $A$ is commutative, $A$ has no $q^{\prime}$-Heisenberg normal element.
For $\mathbb{Z}$-graded $k$-algebra AA, in some situations, the regularity of a $q^{\prime}$-Heisenberg normal element of $A$ is not necessarily, but $u\in k^{\times}$ is sufficient. For convenience, we assume the regularity of a $q^{\prime}$-Heisenberg normal element. We show some examples of a $q^{\prime}$-Heisenberg normal element of some algebras.
Example 3.3.
  1. Suppose that $A(\alpha,\beta)$ is a graded down-up algebra
    $A(\alpha,\beta):=k\langle x,y\rangle/\langle x^{2}y-\alpha xyx-\beta yx^{2},xy^{2}-\alpha yxy-\beta y^{2}x\rangle.$
    where $\alpha,\beta\in k$ (see [12]). When $\alpha,\beta\neq 0$ and $\alpha^{2}+4\beta=0$ hold, $g=xy-\dfrac{\alpha}{2}yx$ is a $q^{\prime}$-Heisenberg normal element of $A$. This algebra is isomorphic to another $q$-analog of the universal enveloping algebra of the Heisenberg algebra $H^{\prime}q$ with $q=\dfrac{\alpha}{2}$ as in [13, Proposition 2.4].
  2. (1)
    Other example of a $q^{\prime}$-Heisenberg normal element can be found in [16, Theorem A (d)]. Let $D(v,p)$ be the algebra defined as follows:
    $D(v,p):=k\langle x,y\rangle\left/\middle\langle\begin{aligned} &xy^{2}+vyxy+p^{2}y^{2}x,\\ &x^{3}y+(v+p)x^{2}yx+(p^{2}+pv)xyx^{2}+p^{3}yx^{3}\end{aligned}\right\rangle$
    where $v,p\in k$ and $p\neq 0$. When $v=2p$ holds, an element $x^{2}y+2pxyx+p^{2}yx^{2}\in D(v,p)$ is a $q^{\prime}$-Heisenberg normal element.
  3. (2)
    More generally, for a $\mathbb{N}$-graded $k$-algebra AA, let $f$ be a homogeneous element with degree $n$, and $\delta$ a graded derivation of $A$. Suppose that $B=A[X,\delta]$ is a graded Ore extension of $A$ with a degree $1$ element $X$. If $Y=\delta(f)$ is regular in $B[Y]/\langle Y-\delta(f)\rangle$ where $Y$ is a degree $n+1$ element, then $\delta(f)$ is a $q^{\prime}$-Heisenberg normal element of it. Indeed, the definition of the Ore extension gives the following equations;
    $\begin{cases}Xf=fX+\delta(f)&\text{ in }A[X,\delta];\\ Yr=rY&\text{ for }r\in B[Y]\text{ in }B[Y]=A[X,\delta][Y].\end{cases}$
We will show some basic properties about a $q^{\prime}$-Heisenberg normal element. First, we recall a basic property of a normal element.
Lemma 3.4.
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra, $g$ a regular homogeneous normal element of $A$ with degree $n\geq 1$, and $A^{[n]}$ the $n$-th quasi-Veronese algebra of $A$. Suppose that $\mathbf{g}$ is the element of $A^{[n]}$ defined as
$\mathbf{g}:=(\delta_{i,j}g)=\left(\begin{array}[]{cccc}g&&&\\ &g&&\\ &&\ddots&\\ &&&g\\ \end{array}\right)\in\left(\begin{array}[]{cccc}A_{n}&A_{n+1}&\cdots&A_{2n-1}\\ A_{n-1}&A_{n}&&\vdots\\ \vdots&&\ddots&\vdots\\ A_{1}&\cdots&\cdots&A_{n}\\ \end{array}\right)=(A^{[n]})_{1},$
where $\delta_{i,j}$ is the Kronecker delta. Then $\mathbf{g}$ is a regular homogeneous normal element of $A^{[n]}$ with degree $1$.
Proof.
Let $\nu_{g}$ the graded automorphism determined by $\nu_{g}(a)g=ga$ for $a\in A$. Also, the map defined by $\nu_{\mathbf{g}}^{[n]}((a_{i,j}))=(\nu_{g}(a_{i,j}))$ for $(a_{i,j})\in(A^{[n]})_{r}$ is a graded automorphism of $A^{[n]}$ in [18]. For $\mathbf{a}=(a_{i,j})\in(A^{[n]})_{r}$, the following equation holds;
$\mathbf{ga}=\left(\sum_{l=0}^{n-1}\delta_{l,j}ga_{i,l}\right)=(ga_{i,j})=(\nu_{g}(a_{i,j})g)=\left(\sum_{l=0}^{n-1}\nu_{g}(a_{l,j})\delta_{i,l}g\right)=\nu_{\mathbf{g}}^{[n]}(\mathbf{a})\mathbf{g}.$
Then $\mathbf{g}$ is a regular normal element in $A^{[n]}$ with degree $1$. ∎
Note that, by [3] and [31, Example 2.12], the set $\nu_{\mathbf{g}}^{[n]}=\{(\nu_{\mathbf{g}}^{[n]})^{i}\mid i\in\mathbb{Z}\}$ is a twisting system of $A^{[n]}$ and $\mathbf{g}$ is a central element of the twisted algebra $^{\nu_{\mathbf{g}}^{[n]}}A^{[n]}$.
Remark 3.5.
In the rest of this section, we use the presented notations. For a $\mathbb{Z}$-graded $k$-algebra AA, let $g$ be a regular homogeneous normal element of $A$ with degree $n\geq 1$. We denote an element $\mathbf{g}:=(\delta_{i,j}g)$ of $A^{[n]}$ as $\mathbf{g}$ where $\delta_{i,j}$ is the Kronecker delta. Also, $\nu_{g}$ denotes the graded automorphism of $A$ defined by $\nu_{g}(a)g=ga$ for $a\in A$ and $\nu_{\mathbf{g}}^{[n]}$ denotes the graded automorphism of $A^{[n]}$ defined by $\nu_{\mathbf{g}}^{[n]}((a_{i,j}))=(\nu_{g}(a_{i,j}))$ for $(a_{i,j})\in(A^{[n]})_{r}$.
To investigate a relationship between an algebra having a $q^{\prime}$-Heisenberg normal element and the first Weyl algebra, we recall the definition of the first Weyl algebra $W_{1}(k)$ (see [10]);
$W_{1}(k):=k\langle X,Y\rangle/\langle XY-YX-1\rangle.$
Also, we recall a filtration on the first Weyl algebra called the Bernstein filtration $\mathcal{B}=\{\mathcal{B}_{i}\}$ (see [10]);
$\mathcal{B}_{i}=\begin{cases}\left\{\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{|\alpha|+|\beta|\leq i}c_{\alpha,\beta}Y^{\alpha}X^{\beta}\in W_{1}(k)\middle|c_{\alpha,\beta}\in k\right\}&\text{for}\,i\geq 0;\\ 0&\text{for}\,i<0.\end{cases}$
Proposition 3.6.
[…]
Proof.
We denote the identity of $A^{[n]}$ as $\mathbf{1}$. Let $\varphi:W_{1}(k)\rightarrow\mathcal{R}$ be the $k$-algebra homomorphism defined by mapping of generators as follows;
$X\mapsto\left(\begin{array}[]{cccc}&xg&&\\ &&\ddots&\\ &&&xg\\ x&&&\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle\in(A^{[n]})_{1}+\langle\mathbf{1}-\mathbf{g}\rangle=F_{1}\mathcal{R},$
$Y\mapsto\left(\begin{array}[]{cccc}&&&gy\\ y&&&\\ &\ddots&&\\ &&y&\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle\in(A^{[n]})_{1}+\langle\mathbf{1}-\mathbf{g}\rangle=F_{1}\mathcal{R}.$
Since $\nu_{g}(x)=u^{-1}x$, $\nu_{g}(y)=uy$, and $\nu_{g}(g)=g$, the following equations are obtained:
$\varphi(X)\varphi(Y)$ $=\left(\begin{array}[]{cccc}&xg&&\\ &&\ddots&\\ &&&xg\\ x&&&\\ \end{array}\right)\circ\left(\begin{array}[]{cccc}&&&gy\\ y&&&\\ &\ddots&&\\ &&y&\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle$
$=u^{-1}\left(\begin{array}[]{cccc}&xg&&\\ &&\ddots&\\ &&&xg\\ x&&&\\ \end{array}\right)\left(\begin{array}[]{cccc}&&&gy\\ y&&&\\ &\ddots&&\\ &&y&\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle$
$=u^{-1}\left(\begin{array}[]{cccc}xgy&&&\\ &xgy&&\\ &&\ddots&\\ &&&xgy\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle$
$=\mathbf{g}\left(\begin{array}[]{cccc}xy&&&\\ &xy&&\\ &&\ddots&\\ &&&xy\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle$
and
$\varphi(Y)\varphi(X)$ $=\left(\begin{array}[]{cccc}&&&gy\\ y&&&\\ &\ddots&&\\ &&y&\\ \end{array}\right)\circ\left(\begin{array}[]{cccc}&xg&&\\ &&\ddots&\\ &&&xg\\ x&&&\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle$
$=u\left(\begin{array}[]{cccc}&&&gy\\ y&&&\\ &\ddots&&\\ &&y&\\ \end{array}\right)\left(\begin{array}[]{cccc}&xg&&\\ &&\ddots&\\ &&&xg\\ x&&&\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle$
$=u\left(\begin{array}[]{cccc}yxg&&&\\ &\ddots&&\\ &&yxg&\\ &&&gyx\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle$
$=u\mathbf{g}\left(\begin{array}[]{cccc}yx&&&\\ &\ddots&&\\ &&yx&\\ &&&yx\\ \end{array}\right)+\langle\mathbf{1}-\mathbf{g}\rangle.$
Consequently, it follows that
$\varphi(XY-YX-1)=\varphi(X)\varphi(Y)-\varphi(Y)\varphi(X)-\mathbf{1}$
$\quad=\mathbf{g}\left(\begin{array}[]{cccc}xy&&&\\ &\ddots&&\\ &&xy&\\ &&&xy\\ \end{array}\right)-u\mathbf{g}\left(\begin{array}[]{cccc}yx&&&\\ &yx&&\\ &&\ddots&\\ &&&yx\\ \end{array}\right)-\mathbf{1}+\langle\mathbf{1}-\mathbf{g}\rangle=0.$
Clearly, the homomorphism $\varphi$ preserves the filtrations $\mathcal{B}$ and $F\mathcal{R}$. ∎
Since the first Weyl algebra $W_{1}(k)$ is simple (see [10, 2.2.1 Theorem]), $\mathcal{R}$ has a subalgebra isomorphic to $W_{1}(k)$ by Proposition 3.6.
Proposition 3.7.
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, g=xyuyxg=xy-uyx a $q^{\prime}$-Heisenberg normal element of $A$ with degree $n$. Then there exists a one-to-one correspondence between the set of isomorphism classes of base-point modules over $A$ and $A/\langle g\rangle$.
Proof.
[…]
$\operatorname{dim}_{k}F_{s}\mathcal{N}=\operatorname{dim}_{k}{}^{\nu_{\mathbf{g}}^{[n]}}(Q(M))_{s}=\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{i=0}^{n-1}\operatorname{dim}_{k}M_{sn-i}=cn\quad\text{ for $s>0$.}$
Hence, there exists a good filtration $\Gamma=\{\Gamma_{i}\}_{i\in\mathbb{Z}}$ of $\mathcal{N}$ and exist $c_{1}\neq 0,c_{0}\in\mathbb{Q}$ such that $\operatorname{dim}_{k}\Gamma_{s}=c_{1}s+c_{0}$ for $s\gg 0$. Furthermore, there exists $s_{0}\in\mathbb{Z}$ such that $\Gamma_{i}\subset F_{i+s_{0}}\mathcal{N}$ for $i\geq 0$ by [10, 8.3.2]. Hence, the inequality $cn=\operatorname{dim}_{k}F_{s+s_{0}}\mathcal{N}\geq\operatorname{dim}_{k}\Gamma_{s}=c_{1}s+c_{0}$ holds for $s\gg 0$. This is a contradiction. Therefore, $M$ is not $g$-torsionfree and then $gM=0$ by [7, Lemma 5]. ∎
Remark 3.8.
Let $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ be a $g$-tosionfree left bounded graded module such that $\operatorname{dim}_{k}M_{s}=c_{1}s+c_{0}$ for $s\gg 0$ where $c_{1}\neq 0$, $c_{0}\in\mathbb{Q}$ (e.g., $M$ is a line module). By the similar proof of Proposition 3.7, the dehomogenization of $^{\nu_{\mathbf{g}}^{[n]}}Q(M)$ is a holonomic module of $W_{1}(k)$.
4. $q^{\prime}$-Heisenberg normal elements and point modules
In this section, for a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, we show more details of a relationship between a point scheme of $A$ and a $q^{\prime}$-Heisenberg normal element of $A$. In the end of this section, we show the structure of the set of closed points of the point scheme of the universal enveloping algebra of a color Lie algebra.
4.1. $q^{\prime}$-Heisenberg normal elements and truncated point modules
Throughout this subsection, let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ unless otherwise specified. In this subsection, we show a relationship between a $q^{\prime}$-Heisenberg normal element $g$ of $A$ and a truncated $g$-torsionfree point module over $A$.
First, we recall useful equations obtained by the definition of a $q^{\prime}$-Heisenberg normal element of $A$. For a $q^{\prime}$-Heisenberg normal element g=xyuyxg=xy-uyx of $A$, the equations $x^{2}y-2uxyx+u^{2}yx^{2}=0$ and $xy^{2}-2uyxy+u^{2}y^{2}x=0$ holds. Then the same equations in a down-up algebra holds as in [12].
Lemma 4.1.
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra (not necessarily connected, $\mathbb{N}$-graded, finitely presented, or generated in degree $1$) and g=xyuyxg=xy-uyx a $q^{\prime}$-Heisenberg normal element of $A$ with degree $n$. Then, for an integer $r\geq 1$, the following equations hold;
  1. (1)
    $x^{r}y=ru^{r-1}xyx^{r-1}-(r-1)u^{r}yx^{r}$;
  2. (2)
    $yx^{r}=ru^{-(r-1)}x^{r-1}yx-(r-1)u^{-r}x^{r}y$.
Proof.
See [12, Lemma 2.3]. ∎
The display of a $q^{\prime}$-Heisenberg normal element $g$ of $A$ is not necessarily unique, that is, we may have displays $g=xy-uyx=x^{\prime}y^{\prime}-u^{\prime}y^{\prime}x^{\prime}$ where $x,x^{\prime}\in A_{1}$, $y,y^{\prime}\in A_{n-1}$, and $u,u^{\prime}\in k^{\times}$. However, when $g$ is displayed by g=xyuyxg=xy-uyx, a $g$-torsionfree truncated point module over $A$ has a nice relationship to $x$.
Proposition 4.2.
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, g=xyuyxg=xy-uyx a $q^{\prime}$-Heisenberg normal element of $A$ with degree $n$ of $A$, and $P=\bigoplus_{i=0}^{d}km_{i}$ a truncated $g$-torsionfree point module of length $d+1$ over $A$ where $d\geq 2n-1$. Then $xm_{i}\neq 0$ holds for $0\leq i\leq d-1$.
Proof.
Let $\{p_{i+1}\}_{i=0}^{d-1}$ (resp. $\{q_{i+(n-1)}\}_{i=0}^{d-(n-1)}$, $\{\lambda_{i+n}\}_{i=0}^{d-n}$) be a sequence defined by $p_{i+1}m_{i+1}=xm_{i}$ (resp. $q_{i+(n-1)}m_{i+(n-1)}=ym_{i}$, $\lambda_{i+n}m_{i+n}=gm_{i}$). For $1\leq i\leq d-n$, the following equation holds;
$p_{i+n}\lambda_{i+n-1}m_{i+n}=xgm_{i-1}=ugxm_{i-1}=u\lambda_{i+n}p_{i}m_{i+n}.$
Since $P$ is a truncated $g$-torsionfree point module over $A$, $\lambda_{i+n-1}\neq 0$ and $\lambda_{i+n}\neq 0$. Then the equation $p_{i+n}=0$ holds if and only if $p_{i}=0$ holds for $1\leq i\leq d-n$.
Let $B$ be the set defined as $B=\{i\mid p_{i}=0\text{ for }1\leq i\leq n\}$. We will show that $B=\emptyset$. First, assume that the number of elements in $B$ is $2$ or more. Let $(i,j)$ be a pair of elements in $B$ such that $j>i$ and minimalize jij-i.
  1. (1)
    The case that ji=1j-i=1 holds. A contradiction is obtained from the following equation;
    $0\neq gm_{i}=(xy-uyx)m_{i}=(p_{i+n}q_{i+n-1}-uq_{i+n}p_{i+1})m_{i+n}=0.$
  2. (2)
    The case $j-i\geq 2$ that holds. Lemma 4.1 provides
    $0$ $=q_{j+n-1}p_{j}p_{j-1}\cdots p_{i+1}m_{j+n-1}$
    $=yx^{j-i}m_{i}$
    $=((j-i)u^{-(j-i-1)}x^{j-i-1}yx-(j-i-1)u^{-(j-i)}x^{j-i}y)m_{i}$
    $=((j-i)u^{-(j-i-1)}p_{j+n-1}\cdots p_{i+n+1}q_{i+n}p_{i+1}$
    $\quad-(j-i-1)u^{-(j-i)}p_{j+n-1}\cdots p_{i+n+1}p_{i+n}q_{i+n-1})m_{j+n-1}$
    $=(j-i)u^{-(j-i-1)}p_{j+n-1}\cdots p_{i+n+1}q_{i+n}p_{i+1}m_{j+n-1}.$
    It follows that $q_{i+n}=0$ by the choice of $(i,j)$ and the following equation yields a contradiction;
    $0\neq gm_{i}=(xy-uyx)m_{i}=(p_{i+n}q_{i+n-1}-uq_{i+n}p_{i+1})m_{i+n}=0.$
Therefore, the number of elements in $B$ is less than $2$. Assume that $B=\{i\}$ where $1\leq i\leq n$. If $i=n$, then the following equation holds;
$0$ $=q_{2n-1}p_{n}\cdots p_{1}m_{2n-1}$
$=yx^{n}m_{0}$
$=(nu^{-(n-1)}x^{n-1}yx-(n-1)u^{-n}x^{n}y)m_{0}$
$=(nu^{-(n-1)}p_{2n-1}\cdots p_{n+1}q_{n}p_{1}-(n-1)u^{-n}p_{2n-1}\cdots p_{n+1}p_{n}q_{n-1})m_{2n-1}$
$=nu^{-(n-1)}p_{2n-1}\cdots p_{n+1}q_{n}p_{1}m_{2n-1}.$
It follows that $q_{n}=0$ and a contradiction is obtained from the following equation;
$0\neq gm_{0}=(xy-uyx)m_{0}=(p_{n}q_{n-1}-uq_{n}p_{1})m_{n}=0.$
Hence, $i\neq n$, that is, $p_{n}\neq 0$. When $i\neq n-1$, Lemma 4.1 provides
(4.1) $0$ $=p_{2n-1}\cdots p_{i+n}q_{i+n-1}m_{2n-1}$
$=x^{n-i}ym_{i}$
$=((n-i)u^{n-i-1}xyx^{n-i-1}-(n-i-1)u^{n-i}yx^{n-i})m_{i}$
$=((n-i)u^{n-i-1}p_{2n-1}q_{2n-2}p_{n-1}\cdots p_{i+1}$
$\quad-(n-i-1)u^{n-i}q_{2n-1}p_{n}p_{n-1}\cdots p_{i+1})m_{2n-1}.$
It follows that
$(n-i)p_{2n-1}q_{2n-2}q_{n-1}-(n-i-1)uq_{2n-1}p_{n}q_{n-1}=0.$
When i=n1i=n-1, the same equation also holds. Similarly, the following equation is obtained;
$iu^{2}q_{2n-1}q_{n}p_{1}-(i-1)uq_{2n-1}p_{n}q_{n-1}=0.$
Also, the following equation is obtained;
$0$ $=(xy^{2}-2uyxy+u^{2}y^{2}x)m_{0}$
$=(p_{2n-1}q_{2n-2}q_{n-1}-2uq_{2n-1}p_{n}q_{n-1}+u^{2}q_{2n-1}q_{n}p_{1})m_{2n-1}.$
Consequently, the above equations yield
$\left(\begin{array}[]{ccc}(n-i)&-(n-i-1)&0\\ 0&-(i-1)&i\\ 1&-2&1\\ \end{array}\right)\left(\begin{array}[]{c}p_{2n-1}q_{2n-2}q_{n-1}\\ uq_{2n-1}p_{n}q_{n-1}\\ u^{2}q_{2n-1}q_{n}p_{1}\end{array}\right)=0.$
Since $\mathrm{det}\left(\begin{array}[]{ccc}(n-i)&-(n-i-1)&0\\ 0&-(i-1)&i\\ 1&-2&1\\ \end{array}\right)=n$, $p_{2n-1}q_{2n-2}q_{n-1}$, $q_{2n-1}p_{n}q_{n-1}$, $q_{2n-1}q_{n}p_{1}$ are zero. If $q_{2n-1}\neq 0$, then $p_{n}q_{n-1}=q_{n}p_{1}=0$ and a contradiction is obtained from the following equation;
$0\neq gm_{0}=(xy-uyx)m_{0}=(p_{n}q_{n-1}-uq_{n}p_{1})m_{n}=0.$
Therefore, $q_{2n-1}=0$. When $i\neq n-1$, the equation (4.1) gives the following equation:
$0$ $=(n-i)u^{n-i-1}p_{2n-1}q_{2n-2}p_{n-1}\cdots p_{i+1}-(n-i-1)u^{n-i}q_{2n-1}p_{n}p_{n-1}\cdots p_{i+1}$
$=(n-i)u^{n-i-1}p_{2n-1}q_{2n-2}p_{n-1}\cdots p_{i+1}.$
It follows that $p_{2n-1}q_{2n-2}=0$. When i=n1i=n-1, the same equation also holds. Then
$0\neq gm_{n-1}=(xy-uyx)m_{n-1}=(p_{2n-1}q_{2n-2}-uq_{2n-1}p_{n})m_{2n-1}=0.$
This is contradiction. Hence, $B$ is empty set, that is, $xm_{i}\neq 0$ for $0\leq i\leq d-1$. ∎
Proposition 4.2 shows the nonexistence of a long truncated $g$-torsionfree point module over $A$ where $g$ is a $q^{\prime}$-Heisenberg normal element of $A$.
Theorem 4.3.
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and g=xyuyxg=xy-uyx a $q^{\prime}$-Heisenberg normal element of $A$ with degree $n$. For $d\geq 2n-1$, $A$ has no a truncated $g$-torsionfree point module of length $d+1$ over $A$.
Proof.
The notation in the proof of Proposition 4.2 is retained. Let $\{r_{i}\}_{i=1}^{n+1}$ be a sequence defined by $r_{i}=\dfrac{q_{i+n-2}}{p_{i+n-2}p_{i+n-3}\cdots p_{i}}$ and $P=\bigoplus_{i=0}^{d}km_{i}$ a truncated $g$-torsionfree point module of length $d+1$ over $A$ where $d\geq 2n-1$. The definition of a $q^{\prime}$-Heisenberg normal element of $A$ gives the following equations;
$0=(x^{n-1}g-u^{n-1}gx^{n-1})m_{0}=(p_{2n-1}\cdots p_{n+1}\lambda_{n}-u^{n-1}\lambda_{2n-1}p_{n-1}\cdots p_{1})m_{2n-1}$
and
$0=(yg-u^{-1}gy)m_{0}=(q_{2n-1}\lambda_{n}-u^{-1}\lambda_{2n-1}q_{n-1})m_{2n-1}.$
Then the equation $r_{n+1}=u^{-n}r_{1}$ holds. On the other hand, for $1\leq i\leq n-1$,
$0$ $=p^{-1}_{i+n}p^{-1}_{i+n-1}\cdots p^{-1}_{i}(x^{2}y-2uxyx+u^{2}yx^{2})m_{i-1}$
$=p^{-1}_{i+n}p^{-1}_{i+n-1}\cdots p^{-1}_{i}(p_{i+n}p_{i+n-1}q_{i+n-2}-2up_{i+n}q_{i+n-1}p_{i}+u^{2}q_{i+n}p_{i+1}p_{i})m_{i+n}$
$=(r_{i}-2ur_{i+1}+u^{2}r_{i+2})m_{i+n}.$
It follows that $r_{i+1}-u^{-1}r_{i}=u^{1-i}(r_{2}-u^{-1}r_{1})$ for $1\leq i\leq n$ and then
$0$ $=r_{n+1}-u^{-n}r_{1}=\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}^{n}_{i=1}u^{i-n}(r_{i+1}-u^{-1}r_{i})$
$=\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}^{n}_{i=1}u^{i-n}u^{1-i}(r_{2}-u^{-1}r_{1})=nu^{1-n}(r_{2}-u^{-1}r_{1}).$
Consequently, the following equation leads to contradiction;
$0\neq gm_{0}=(p_{n}q_{n-1}-uq_{n}p_{1})m_{n}=p_{1}p_{2}\cdots p_{n}(r_{1}-ur_{2})m_{n}=0.$
Remark that, since a truncated $g$-torsionfree point module over $A$ is not a $g$-torsionfree module over $A$, Theorem 4.3 is not a corollary of Proposition 3.7.
4.2. Point modules over the universal enveloping algebras of color Lie algebras
Throughout this subsection, for integers $m\geq 1$ and $m\geq i\geq 0$, let $e_{i}\in\mathbb{Z}^{(m+1)}$ denotes $(\delta_{i,0},\delta_{i,1},\cdots,\delta_{i,m})$ where $\delta_{i,j}$ is the Kronecker delta. Also, for a finite dimensional $(\mathbb{Z}^{m+1},\varepsilon)$-color Lie algebra $L$ with $L_{-}=0$, let $L_{\mathbf{1}}$ denotes $\bigoplus_{0\leq i\leq m}L_{e_{i}}$ and $n_{L}$ denotes $n_{L}=\mathrm{max}\{j\mid L_{\mathbf{1}}^{j}\neq 0\}$ where $L_{\mathbf{1}}^{1}=L_{\mathbf{1}}$ and $L_{\mathbf{1}}^{j+1}=[L_{\mathbf{1}}^{j},L_{\mathbf{1}}]$ for $j\geq 1$. Furthermore, the map $|\quad|_{\mathbb{Z}}:h(L)\setminus\{0\}\rightarrow\mathbb{Z}$ is defined as $L_{(a_{0},\cdots,a_{m})}\ni x\mapsto a_{0}+\cdots a_{m}\in\mathbb{Z}$. When $U(L)$ is considered as a $\mathbb{Z}$-graded $k$-algebra with the way of Remark 2.18(2), let $\deg x\in\mathbb{Z}$ denotes the degree of $x\in h(U(L))$.
In this subsection, we use Theorem 4.3 to the universal enveloping algebra of a color Lie algebra. First, we recall the definition of a skew polynomial algebra on $(m+1)$-variables (see [6, 15, 30, 17] for details);
$S_{\omega}=k\langle x_{0},x_{1},\cdots,x_{m}\rangle/\langle x_{i}x_{j}-\omega_{i,j}x_{j}x_{i}\mid 0\leq i,j\leq m\rangle$
where $\omega=(\omega_{i,j})$ is a $(m+1)\times(m+1)$ matrix over $k$ such that $\omega_{i,j}\omega_{j,i}=1$ and $\omega_{l,l}=1$ for $0\leq i,j,l\leq m$. In the following, by assuming that the degrees of $x_{i}$ are $1$, we consider $S_{\omega}$ as $\mathbb{Z}$-graded $k$-algebra generated in degree $1$. We recall an important result of the set of isomorphism classes of point modules over a skew polynomial algebra.
Theorem 4.4 ([6, Theorem 1], [30]).
Let $S_{\omega}$ be a skew polynomial algebra on $(m+1)$-variables with $\omega=(\omega_{i,j})$ for $m\geq 2$. Then there exists a one-to-one correspondence between the set of isomorphism classes of point modules over $S_{\omega}$ and
$\{(p_{0}:p_{1}:\cdots:p_{m})\in\mathbb{P}^{m}\mid(\omega_{i,j}\omega_{j,l}-\omega_{i,l})p_{i}p_{j}p_{l}=0\text{ for }0\leq i<j<l\leq m\}.$
Also, when $m=1$, there exists a one-to-one correspondence between the set of isomorphism classes of point modules over $S_{\omega}$ and the set of closed points of $\mathbb{P}^{1}$ by [2]. By those facts and Theorem 4.3, we obtain the structure of the set of isomorphism classes of point modules over the universal enveloping algebra of color Lie algebras. Before starting a main theorem, we remark that, for a finite dimensional $(\mathbb{Z}^{m+1},\varepsilon)$-color Lie algebra $L$ with $L_{-}=0$ generated by $\{\theta_{0},\theta_{1},\cdots,\theta_{m}\}$ with $|\theta_{i}|=e_{i}$ for $0\leq i\leq m$, $S_{\varepsilon}(L_{\mathbf{1}})$ is a skew polynomial algebra since $S_{\varepsilon}(L_{\mathbf{1}})\cong S_{\varepsilon}(L/\bigoplus_{|a|_{\mathbb{Z}}\geq 2}L_{a})$.
Theorem 4.5.
Let $L$ be a finite dimensional $(\mathbb{Z}^{m+1},\varepsilon)$-color Lie algebra with $L_{-}=0$ generated by $\{\theta_{0},\cdots,\theta_{m}\}$ with $|\theta_{i}|=e_{i}$ for $0\leq i\leq m$. Then, for $d\geq 2n_{L}-1$, there exists a one-to-one correspondence between the isomorphism classes of the set of truncated point modules of length $d+1$ over $U(L)$ and the one of $S_{\varepsilon}(L_{\mathbf{1}})$. In particular, for $m\geq 2$, the set of isomorphism classes of point modules over $U(L)$ corresponds one-to-one to
$\{(p_{0}:p_{1}:\cdots:p_{m})\in\mathbb{P}^{m}\mid(\omega_{i,j}\omega_{j,l}-\omega_{i,l})p_{i}p_{j}p_{l}=0\text{ for }0\leq i<j<l\leq m\}$
where $\omega_{i,j}:=\varepsilon(e_{i},e_{j})$ for $0\leq i,j\leq m$. Also, when $m=1$, the set of isomorphism classes of point modules over $U(L)$ corresponds one-to-one to the set of closed points of $\mathbb{P}^{1}$;
Proof.
When $n_{L}=1$, since $L=L_{\mathbf{1}}$, the argument is clear. We use induction on $\operatorname{dim}_{k}L$. If $\operatorname{dim}_{k}L=2$ holds, then $n_{L}=1$ holds. Suppose that $\operatorname{dim}_{k}L\geq 3$ and $n_{L}\geq 2$. Let $g=[x,y]\neq 0$ be an element of $L_{\mathbf{1}}^{n_{L}}$. By reselecting $g$ if necessary, one can assume that $g=[x,y]\neq 0$ satisfies $x\in h(L_{\mathbf{1}})$ and $y\in h(L)$. Since $L$ is generated by $L_{\mathbf{1}}$, $[g,L]=0$ holds. Then $g$ is a normal element of $U(L)$. Also, $g$ is regular since $U(L)$ is a domain([21]). Furthermore, the equations $\varepsilon(|x|,|x|+|y|)=\varepsilon(|x|,|x|)\varepsilon(|x|,|y|)=\varepsilon(|x|,|y|)$ and $\varepsilon(|x|+|y|,|y|)=\varepsilon(|x|,|y|)\varepsilon(|y|,|y|)=\varepsilon(|x|,|y|)$ hold. It follows that
[…]
Hence, $g$ is a $q^{\prime}$-Heisenberg normal element of $U(L)$ with $\deg g=n_{L}\geq 2$. Therefore, for $d\geq 2n_{L}-1$, the set of isomorphism classes of truncated point modules of length $d+1$ over $U(L)$ corresponds one-to-one to the one of $U(L)/\langle g\rangle$ by Proposition 2.13 and Theorem 4.3. By the similar proof of [11, 2.2.14], we have $U(L)/\langle g\rangle\cong U(L/\langle g\rangle)$. Also, clearly $d\geq 2n_{L}-1\geq 2n_{L/\langle g\rangle}-1$ and $S_{\varepsilon}(L_{\mathbf{1}})\cong S_{\varepsilon}((L/\langle g\rangle)_{\mathbf{1}})$. By induction hypothesis, we have conclusion. The second claim is follows from Theorem 4.4. ∎
Note that the second claim of Theorem 4.5 can also be followed by Proposition 3.7. On the other hands, the following corollary is only followed by Theorem 4.3.
Corollary 4.6.
Let $L$ be a finite dimensional $(\mathbb{Z}^{m+1},\varepsilon)$-color Lie algebra with $L_{-}=0$ and $n_{L}\geq 2$ generated by $\{\theta_{0},\cdots,\theta_{m}\}$ with $|\theta_{i}|=e_{i}$ for $0\leq i\leq m$. Suppose that $\Gamma_{j}(U(L))$ is the $j$-th truncated point scheme of $U(L)$ and $\Gamma(U(L))$ is the point scheme of $U(L)$. Then $\mathrm{pr}_{1,2n_{L}-2}^{(2n_{L}-1)}\mathrm{:}\,\Gamma_{2n_{L}-1}(U(L))\rightarrow\Gamma_{2n_{L}-2}(U(L))$ is a closed immersion and identify a closed subscheme $E\subset\Gamma_{2n_{L}-2}(U(L))$. Moreover, $\Gamma(U(L))\cong E$ as schemes.
Proof.
For a scheme $X$, the set of closed points of $X$ is denoted by $\mathrm{Cl}(X)$. By Theorem 4.5, we obtain an equation $\mathrm{Cl}(\Gamma_{2n_{L}-1}(U(L)))=\mathrm{Cl}(\Gamma_{2n_{L}-1}(S_{\varepsilon}(L_{\mathbf{1}})))$. Also, for a normal element $g$ of $U(L)$ where $g\in h(L^{n_{L}}_{\mathbf{1}})$, Propositions 2.13 and 2.14 provide $\mathrm{Cl}(\Gamma_{2n_{L}-2}(U(L/\langle g\rangle)))\subset\mathrm{Cl}(\Gamma_{2n_{L}-2}(U(L)))$. Inductively, we obtain
$\mathrm{Cl}(\Gamma_{2n_{L}-2}(S_{\varepsilon}(L_{\mathbf{1}})))\subset\mathrm{Cl}(\Gamma_{2n_{L}-2}(U(L))).$
Then, by [25, 11.2.Q.Exercise] and [6] (also, see [30]), $\mathrm{pr}_{1,2n_{L}-2}^{(2n_{L}-1)}:\Gamma_{2n_{L}-1}(U(L))\rightarrow\Gamma_{2n_{L}-2}(U(L))$ is injective on points. By [2, Proposition 3.6(ii)], $\mathrm{pr}_{1,2n_{L}-2}^{(2n_{L}-1)}$ is a closed immersion and identify a closed subscheme $E\subset\Gamma_{2n_{L}-2}(U(L))$. Also, $\sigma:E\rightarrow\Gamma_{2n_{L}-2}(U(L))$ is defined as $(p_{1},\cdots,p_{2n_{L}-2})\mapsto(p_{2},\cdots,p_{2n_{L}-1})$ where $(p_{1},\cdots,p_{2n_{L}-1})\in\Gamma_{2n_{L}-1}(U(L))$ is the point laying over $(p_{1},\cdots,p_{2n_{L}-2})$ by [2, Proposition 3.7(i)]. By similar discussion in above, $\sigma(E)\subset E$ is obtained.
Let $I$ be an ideal of $k\langle x_{0},\cdots x_{m}\rangle$ such that $U(L)\cong k\langle x_{0},\cdots x_{m}\rangle/I$ where the $\deg x_{j}=1$ for $0\leq j\leq m$ and $B=\{b_{1},\cdots,b_{\operatorname{dim}_{k}L}\}$ a homogeneous $k$-vector basis of $L$. We denote the subset $B_{2}$ of $B\times B$ defined as $B_{2}=\{(b_{\ell},b_{\ell^{\prime}})\in B\times B\mid 1\leq\ell<\ell^{\prime}\leq\operatorname{dim}_{k}L\}$. Suppose that $C_{\bullet}=(C_{i},d_{i})$ is the color Koszul resolution of $U(L)$ in Theorem 2.19 and $P_{\bullet}=(P_{i},\partial_{i})$ is a minimal free resolution of $k$ as $\mathbb{Z}$-graded $U(L)$-module. There exist chain maps $f_{\bullet}:P_{\bullet}\rightarrow C_{\bullet}$ and $\pi:C_{\bullet}\rightarrow P_{\bullet}$ such that $\pi f=\mathrm{id}$. Clearly, $P_{2}$ is the summand of $C_{2}\cong\operatorname*{\mathchoice{\Xdbigoplus}{\Xdbigoplus}{\Xtbigoplus}{\Xtbigoplus}}\limits_{(b,b^{\prime})\in B_{2}}U(L)(-|b|_{\mathbb{Z}}-|b^{\prime}|_{\mathbb{Z}})$. Assume that $P_{2}$ has a direct sum decomposition $P_{2}=Q\oplus Q^{\prime}$ such that $Q\cong U(L)(-2n_{L})$. Let $w$ be a free basis of $Q$ and $f_{2}(w)=\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{(b,b^{\prime})\in B_{2}}v_{(b,b^{\prime})}\otimes(b\wedge b^{\prime})$ the image of $w$ where $v_{(b,b^{\prime})}\in h(U(L))$ with $\deg v_{b,b^{\prime}}+|b|_{\mathbb{Z}}+|b^{\prime}|_{\mathbb{Z}}=2n_{L}$. Suppose that $z$ is an element of $C_{2}$ defined as $z=\operatorname*{\mathchoice{\Xdsum}{\Xdsum}{\Xtsum}{\Xtsum}}_{|b|_{\mathbb{Z}}+|b^{\prime}|_{\mathbb{Z}}<2n_{L}}v_{(b,b^{\prime})}\otimes(b\wedge b^{\prime})$. If $|b|_{\mathbb{Z}}+|b^{\prime}|_{\mathbb{Z}}<2n_{L}$, then $\pi_{2}(1\otimes(b\wedge b^{\prime}))\in Q^{\prime}$. It follows that $\pi_{2}(z)\in Q^{\prime}$. On the other hand, if $|b|_{\mathbb{Z}}+|b^{\prime}|_{\mathbb{Z}}=2n_{L}$, then $|b|_{\mathbb{Z}}=|b^{\prime}|_{\mathbb{Z}}=n_{L}$. Since $U(L)$ is generated by degree $1$ and $n_{L}\geq 2$, the following equation holds;
$\partial_{2}\pi_{2}(v_{(b,b^{\prime})}\otimes(b\wedge b^{\prime}))$ $=\pi_{1}d_{2}(v_{(b,b^{\prime})}\otimes(b\wedge b^{\prime}))$
$=\pi_{1}\left(v_{(b,b^{\prime})}b\otimes b^{\prime}-\varepsilon(|b|,|b^{\prime}|)v_{(b,b^{\prime})}b^{\prime}\otimes b\right)=0.$
It follows that $\partial_{2}(w)=\partial_{2}\pi_{2}f_{2}(w)=\partial_{2}\pi_{2}(z)$ and then $\partial_{2}(P_{2})\subset\partial_{2}(Q^{\prime})$ is obtained. This contradicts the minimality of $P_{\bullet}$. Hence, $P_{2}$ is isomorphic to a direct summand of $\operatorname*{\mathchoice{\Xdbigoplus}{\Xdbigoplus}{\Xtbigoplus}{\Xtbigoplus}}\limits_{|b|_{\mathbb{Z}}+|b^{\prime}|_{\mathbb{Z}}<2n_{L},\,(b,b^{\prime})\in B_{2}}U(L)(-|b|_{\mathbb{Z}}-|b^{\prime}|_{\mathbb{Z}})$. This means that $I$ is generated by relations with degree $\leq 2n_{L}-1$ or less by [23, Lemma 2.1.3]. Therefore, by [2, Proposition 3.7(ii)], for $r\geq 2n_{L}$, $\mathrm{pr}_{1,2n_{L}-1}^{(r)}:\Gamma_{r}(U(L))\rightarrow E$ is isomorphism and then $\Gamma_{U(L)}=\displaystyle{\lim_{\leftarrow}}\Gamma_{j}(U(L))\cong E$. ∎
Acknowledgments
I am deeply grateful to my Ph.D. supervisor, Professor Ayako Itaba, for her invaluable guidance and useful discussions. I would like to thank Shota Inoue for his helpful comments on this research.